Circular No. 7185

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999cl IN NGC 4501 M. Papenkova, A. V. Filippenko, and R. R. Treffers, University of

California at Berkeley, on behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report their discovery of an apparent supernova in an unfiltered image taken on May 29 UT (mag 16.4) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). The object was confirmed in images taken on May 30 and 31 (mag 16.2 and 15.3, respectively). Images from May 21 and 22 do not show the object, to a limiting magnitude of 18.6. The supernova candidate is at α = 12^{h}31^{m}55^{s}.9, δ = +14^{o}25′33′′ (equinox 2000.0), which is ∼ 46′′ west and 23′′ north of the nucleus of NGC 4501.

V382 VELORUM J. E. Steiner and R. Campos, Laboratorio Nacional de Astrofisica

(LNA), CNPq; and D. Cieslinski, Instituto Nacional de Pesquisas Espaciais, write: “We observed V382 Vel with the LNA 1.6-m telescope (+ Coudé spectrograph) on May 24. Three exposures were taken, centered at Hα, Hβ, and Ca II H and K. The spectrum shows very strong Fe II emission lines at 492.4 and 501.9 nm (multiplet 42) with about half the intensity of Hβ. Hα presents an equivalent width of −26.3 nm, FWHM of 2390 km/s, and FWZI of 6800 km/s. The interstellar Ca II K line has an equivalent width of 0.024 nm. The derived distance is 800 pc. Assuming a reddening of A_{v} = 0.8 mag, the absolute magnitude at maximum light is M_{v} = −7.7. We classify the nova as of ‘Fe II’ type in the Williams classification system. With the observed ejection velocity, it is likely to be a neon nova. If the pre-outburst magnitude is 16.4 (IAUC 7177), the amplitude of the eruption is 13.8 and the pre-outburst absolute magnitude is M_{v} = +6.1.”

J. Garcia, San Rafael, Mendoza, Argentina, reports that he visually scanned the region surrounding this nova with a telescope in the course of his observation of FW Vel on May 21.083 UT, and he saw nothing unusual in the vicinity of the nova to an estimated magnitude of 10.5.

SUPERNOVAE 1998S, 1999Z, 1999aa, 1999ac, 1999an, 1999be R. Szabo reports CCD R magnitude estimates of various supernovae

using the 1-m telescope at Piszkéstető, Hungary (based on USNO-A1.0 R magnitudes): SN 1998S, May 25.978 UT, 18.4; SN 1999Z, May 25.916, 16.8; SN 1999aa, May 24.820, 18.2; SN 1999ac, May 20.977, 16.1; SN 1999an, May 25.987, 16.5; SN 1999be, May 21.819, 17.8.

1999 June 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green